a. Suppose that a market research company finds that at a price of p = $20, they would sell x = 42 tiles each month. If they lower the price to p = $10, then more people would purchase the tile, and they can expect to sell x = 52 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b. (Hint: Write an equation using two points in the form (x,p)).
Answer by stanbon(75887)

Suppose that a market research company finds that at a price of p = $20, they would sell x = 42 tiles each month. If they lower the price to p = $10, then more people would purchase the tile, and they can expect to sell x = 52 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b.
---------------------------------------
(42,20) and (52,10)
slope = (10-20)/(52-42) = -1
-----------------
Form p = mx + b where p = 20 when x = 42 and m=-1; solve for b:
20 = -1*42 + b
b = 62
-------------
EQUATION:
p = -x + 62
==================
Cheers,
Stan H.